Solve for 7-3(8t-10)= 85
Answer by checkley77(12569)

7-3(8t-10)= 85
7-24t+30=85
-24t=85-30-7
-24t=48
t=48/-24
t=-2 ans.
Proof:
7-3(8*-2-10)=85
7-3(-16-10)=85
7-3*-26=85
7+78=85
85=85